# Everybody Prune Now: Structured Pruning of LLMs with Only Forward Passes

### v3

@@ 4 Analysis
We conduct ablative experiments to understand the impact of the ingredients from Section 2.
Do we need both perturbative and regressive components of Bonsai? Figure 4 shows that both components are key to obtaining a good pruned model. Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl). Further degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform perturbative evaluations on the parent model but simply prune according to the prior $\rho$ as computed from the unperturbed parent model.
@@
Table 5: Wikitext-2 perplexity of LLaMA-2 7B @ 50% sparsity (No PPA). We vary the number of perturbative evaluations. Details in Appendix I.
@@
Table 6: Impact of PPA on LLaMA-2 7B @ 50% sparsity. Details in Appendix G.
@@
Figure 5: LLaMA-2 7B pruned to 50% sparsity. See Appendix H for experiment details and definitions of Wanda and Activation Magnitude priors.
@@
5 Related Work
Here we discuss relevant work in LLM pruning, and defer a longer discussion compression methods to Appendix B.
Unstructured pruning:
Whilst structured pruning approaches remove entire model components like layers (Xu et al., 2020; Xia et al., 2022b), dimensions of linear layers (Wang et al., 2019) or attention heads (Michel et al., 2019; Held & Yang, 2022), unstructured pruning (Han et al., 2015; Frankle & Carbin, 2018; Benbaki et al., 2023; Sun et al., 2023) removes individual parameters of the model. These approaches achieve memory savings by inducing sparsity in the model weights, but they generally do not result in actual model speedups except when specialized hardware is available (Mishra et al., 2021). Proposed semi-structured sparsity methods (Mishra et al., 2021) such as 2:4 and 4:8 patterns do result in faster inference, but the speedup gains they achieve are far from the idealized 2×2$\times$2 ×.
Structured pruning without backward passes:
To the best of our knowledge, all existing structured pruning techniques for large (over 1B scale) language models, like LoRAPrune (Zhang et al., 2023) and LLM-Pruner (Ma et al., 2023), are gradient-based. Thus, under the memory setting we consider, these methods cannot be applied since their memory requirements well exceed the memory needed for inference on the model to be pruned. For smaller language models like BERT, (Nova et al., 2023) recently proposed Kernelized Convex Masking (KCM) for gradient-free structured pruning. Unfortunately, to prune a fully connected layer with $K$ intermediate units, KCM requires instantiating a $K\times K$ coefficient matrix (for each layer). Whilst this is feasible for BERT models, a typical LLM (AI, 2023; Touvron et al., 2023) has $K\sim 10^{4}-10^{5}$ which would make the size of each per-layer coefficient matrix comparable to the size of the LLM itself. In computer vision, there exist several perturbative gradient-free structured pruning techniques (Ancona et al., 2020; Dekhovich et al., 2021). However, these methods have been exclusively applied to small vision models (VGG; Simonyan & Zisserman (2014), WideResnet; Zagoruyko & Komodakis (2016)) and, taken as is, would not scale to LLMs. We introduce several techniques to make Bonsai scale to large language models. A key difference between Bonsai and these perturbative methods is that pruning decisions are made globally instead of on a local layer-by-layer basis (Ancona et al., 2020; Dekhovich et al., 2021). A failure mode of layer-by-layer pruning is that if a module, say $m_{i}$, which is in the top $p$ fraction of relevant modules when the whole model is considered but in the bottom $1-p$ fraction within its layer will not be included in the pruned model — a clearly sub-optimal choice. Bonsai enables us to obtain globally meaningful estimates of module relevance that are comparable across layers.
@@ 6 Conclusion, Limitations, and Future Work
In this work, we have presented Bonsai, a gradient-free structured pruning method that enables efficient compression of LLMs using only forward passes. By eliminating the need for backward passes, Bonsai reduces memory requirements by over 2×2$\times$2 ×, allowing practitioners to prune and adapt models on hardware configurations where gradient-based approaches would be infeasible. Through extensive experiments, we have demonstrated that Bonsai produces models that are small, fast, and accurate—-outperforming even most of these gradient-based methods while using substantially fewer resources.
A primary limitation of Bonsai is its computational runtime. As shown in Section 4, performance improves with more sub-model exploration, slower pruning rates, and larger data samples, but these improvements come at the cost of increased runtime. While competitors like FLAP complete pruning in $\approx 1$ hour, our configuration of Bonsai that achieves optimal results requires $\leq 4$ hours. However, Bonsai offers practitioners the flexibility to configure faster pruning schedules (as quick as 15 minutes) when speed is prioritized over achieving optimal compression quality. This runtime-quality tradeoff is a design feature that allows adaptation to various computational constraints. Additionally, since the models produced by Bonsai can guarantee optimal performance and speedups, this one-time investment becomes worthwhile when amortized over large-scale deployment.
Bonsai presents several promising avenues for future research. First, while our current approach samples sub-models from an informative prior, the sampling process is not completely adaptive. Bonsai could be further enhanced by dynamically exploring the space of sub-models based on previous evaluations. Additionally, due to our focus on forward-pass-only operations, Bonsai does not fine-tune the model during iterative pruning. Integration with gradient-free approaches like MeZO (Malladi et al., 2023) could allow for continuous updates during the pruning process, potentially leading to better preservation of model performance.
Appendix A Broader Impact Statement
@@
Appendix B Background: Additional Methods for LLM Compression
Although we focus on pruning in this work, we note that prior research has explored various other compression schemes such as distillation (Hinton et al., 2015; Gu et al., 2023) and quantization (Xiao et al., 2023) to create smaller models from larger pre-trained ones. Similar to structured pruning, these compression methods themselves often impose significant computational burdens. For example, distillation-based techniques require using LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023). Although unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization have lower training-time resource demands, the models they produce either require specialized hardware to achieve speedups (Mishra et al., 2021) or may actually slow down inference due to additional computational overhead (Dettmers et al., 2022). Table 7 summarizes these existing limitations across different compression approaches. We also cover additional background in distillation and quantization below.
Table 7: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver. ✗  $\rightarrow$ a prohibitive cost to the lay practitioner whilst ✓  $\rightarrow$ a viable option with respect to that resource.
Regime Resource Approaches
Train Memory
Compute
Inference Memory
Compute
Distillation:
Originally proposed by (Hinton et al., 2015; Sanh et al., 2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation. The resulting models are lighter and faster at inference time. Unfortunately, as noted by (Xia et al., 2022b), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance. To reduce this resource burden, (Sanh et al., 2020; Lagunas et al., 2021; Xia et al., 2022b) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch. Bonsai also leverages distillation, but only during the post-pruning adaptation phase. This maintains the constraint that the pruning process is gradient-free. And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
Quantization:
(Li et al., 2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once. (Dettmers et al., 2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition. (Xiao et al., 2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques. This paper focuses on gradient-free structured pruning. We leave combining Bonsai with quantization for future work.
Appendix C Main Experiment Details
C.1 Full Algorithm
Algorithm 1 Bonsai
1:  Input: Model [$\mathbf{M}_{\theta}$], sub-models per iteration [$n_{\mathrm{iter}}$] Sparsity per iteration [$p_{\mathrm{iter}}$], Target sparsity [$p$] Module list [$\mathbf{m}$]
2:  for $l=1$ to $\lceil\frac{p}{p_{\mathrm{iter}}}\rceil$ do
3:     $\mathbb{\rho}^{l}\leftarrow$ Calculate unstructured pruning metric for all modules in $\mathbf{m}$
4:     $\bar{\mathbb{\rho}}^{l}\leftarrow$ Fix the top $(1-2p_{\mathrm{iter}})$ of $\mathbb{\rho}^{l}$ to $\infty$
5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
6:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
7:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
8:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.
9:     $\mathbf{m}\leftarrow$ update module list to exclude $\{m_{\mathrm{pruned}}\}$
10:  end for
11:  Output: Pruned model $\mathbf{M}_{|\mathbf{m}}$
A note about Line $5$ in our algorithm. Depending on the task, we find that sampling $\mathbf{\bar{m}}_{i}$ and its complement $\mathbf{\bar{m}}^{c}_{i}$ helps reduce the variance of our regression estimate and leads to better results.
C.2 Hyper-parameters for all Bonsai regression during pruning
When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma & Ba, 2014). We cross-validate over the following set of hyper-parameters. Note that doing this cross-validation takes much less time than the time needed to construct the dataset $\mathbb{D}^{l}$.
@@ C.2 Hyper-parameters for all Bonsai regression during pruning
During cross validation, we choose the model whose predictions have the best Kendall rank correlation co-efficient (Kendall, 1948) with the target. We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amounts sub-models, which would denote that $\beta$ reasonably models relative module importances.
In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 3.2.3, we find that $\ell_{2}$-norm works better.
C.3 Forward Pass Only / Semi-structured pruning Experiments
Table 10 show the Bonsai hyperparameters we used for the experiments in Section 3.2.1.
@@ C.3 Forward Pass Only / Semi-structured pruning Experiments
C.4 Experiments comparing to Gradient based structured pruning
@@
All Bonsai hyper-parameters are the same as Appendix C.3 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
C.5 Phi-2 pruning experiment details
For the experiment in Section 3.2.3, All Bonsai hyper-parameters are the same as Appendix C.3 except for the following changes:
@@ C.5 Phi-2 pruning experiment details
Appendix D Impact of regression and perturbation ablation details
For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this. This is the No Perturbation + No Regression baseline in Figure 4. This approach has quite poor performance. We can further improve this baseline by adding back perturbative aspect where we prune the parent model according to $\rho^{\prime}$ which is computed by aggregating the $\rho$ metric computed over the perturbed models. Note that we use a Wanda based metric to define $\rho$ for this experiment. Module level analogues of the unstructured pruning metrics we explore are defined in Appendix H.
@@ Appendix D Impact of regression and perturbation ablation details
Linear Regression Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
@@
Appendix E Varying the pruning fraction per-iteration
For the experiment in Section 6, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $p_{\mathrm{iter}}$.
@@ Appendix E Varying the pruning fraction per-iteration
Prune Frac Relative Speepdup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
@@
Appendix F Varying the number of calibration data points for pruning
All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
Table 13: How many data-points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
@@ Appendix F Varying the number of calibration data points for pruning
Appendix G Post-pruning adaptation
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3.
Appendix H Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\rho$.
H.1 $\rho$ is Activation Magnitude
@@
Self-Attention Module: For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho$.
@@
H.2 $\rho$ is Wanda (Sun et al., 2023)
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP. For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module level analogue of Wanda:
@@
Self-Attention Module: Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the self-attention module. For any data-sample sequence $b$, the output of the self attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
H.3 $\rho$ is Fluctuation-based (FLAP) An et al. (2024)
@@
$\left(\rho\in\mathbb{R}^{d}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (7)
where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $||W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
@@
$\left(\rho\in\mathbb{R}^{h}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (8)
@@
Appendix I How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.

### v4

@@ 5 Analysis
Here we conduct ablative experiments to understand the impact of the ingredients from Section 3.
Do we need both perturbative and regressive components of Bonsai?
Figure 3: LLaMA-2 7B @ 50% sparsity (No PPA). Perturbation and regression components are both needed to make Bonsai effective. Experiment details in Appendix B.
Figure 3 shows that both components are key to obtaining a good pruned model. Removing the estimation of module importances via regression leads to a degradation in performance ($61.6$ ppl $\rightarrow$ $146.6$ ppl). Further degradation ($146.6$ ppl $\rightarrow$ $405.7$ ppl) is encountered if we do not perform perturbative evaluations on the parent model but simply prune according to the prior $\rho$ as computed from the unperturbed parent model.
@@
Table 7: Wikitext-2 perplexity of LLaMA-2 7B @ 50% sparsity (No PPA). We vary the number of perturbative evaluations. Details in Appendix F.
@@
Table 8: Impact of PPA on LLaMA-2 7B @ 50% sparsity. Details in Appendix D.
@@
Figure 4: LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details and definitions of Wanda and Activation Magnitude priors.
@@
Memory and Runtime Considerations. Bonsai requires $\approx 20$GB memory versus 80-160GB for gradient-based methods (Table 2). Runtime is $\approx 4$ hours for optimal quality or $\approx 15$ minutes with reduced perturbations. Pre-adaptation results appear in Table 6 and 8; note that gradient-based baselines in Table 4 also use adaptations (Section 4.2).
@@ 6 Conclusion, Limitations, and Future Work
In this work, we have presented Bonsai, a gradient-free structured pruning method that enables efficient compression of LLMs using only forward passes. By eliminating the need for backward passes, Bonsai reduces memory requirements by over $2\times$, allowing practitioners to prune and adapt models on hardware configurations where gradient-based approaches would be infeasible. Through extensive experiments, we have demonstrated that Bonsai produces models that are small, fast, and accurate—outperforming even most of these gradient-based methods while using substantially fewer resources.
A primary limitation of Bonsai is its computational runtime. As shown in Section 5, performance improves with more sub-model exploration, slower pruning rates, and larger data samples, but these improvements come at the cost of increased runtime. While competitors like FLAP complete pruning in $\approx 1$ hour, Bonsai’s optimal configuration requires $\leq 4$ hours. However, Bonsai offers practitioners the flexibility to configure faster pruning schedules ($\approx$15 minutes) when speed is prioritized over achieving optimal compression quality. This runtime-quality tradeoff is a design feature that allows adaptation to various computational constraints. Additionally, since the models produced by Bonsai can guarantee optimal performance and speedups, this one-time investment becomes worthwhile when amortized over large-scale deployment.
Bonsai presents several promising avenues for future research. First, while our current approach samples sub-models from an informative prior, the sampling process is not completely adaptive. Bonsai could be further enhanced by dynamically exploring the space of sub-models based on previous evaluations. Additionally, due to our focus on forward-pass-only operations, Bonsai does not fine-tune the model during iterative pruning. Integration with gradient-free approaches like MeZO (Malladi et al., 2023) could allow for continuous updates during the pruning process, preserving model performance. MoE models also provide an interesting direction: while they enable sparse activation, all experts must remain in memory. Bonsai could prune individual experts to reduce memory footprint while preserving conditional computation benefits.
Overall, by reducing the memory and computational requirements of structured pruning, Bonsai contributes to the growing trend of practical, efficiency-focused AI research, aligning with the community-wide push to make powerful models more accessible. The ability that Bonsai enables to prune and adapt large models on a wider range of hardware configurations fosters innovation beyond resource-rich institutions to represent a crucial move toward a more inclusive, sustainable future for the field.
7 Acknowlegements
This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Meta, Morgan Stanley, Amazon, and Google. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies. We are grateful for helpful feedback from Mingjie Sun, Victor Akinwande, Asher Trockman, Afshin Rostamizadeh and Daniel Glasner.
8 Broader Impact Statement
@@
Appendix A Main Experiment Details
A.1 Hyper-parameters for all Bonsai regression during pruning
When using Bonsai, we estimate $\beta$ from $\mathbb{D}^{l}$ by performing linear regression via gradient descent with Adam (Kingma and Ba, 2014). We cross-validate over the following set of hyper-parameters. Note that doing this cross-validation takes much less time than the time needed to construct the dataset $\mathbb{D}^{l}$.
@@ A.1 Hyper-parameters for all Bonsai regression during pruning
During cross validation, we choose the model whose predictions have the best Kendall rank correlation coefficient (Kendall, 1948) with the target. We do this because we do not care about matching $U_{k}$ exactly for each sub-model $k$; we rather care that our learned $\beta$ predicts the correct rankings amoung sub-models, which would denote that $\beta$ reasonably models relative module importances.
In general, we use $\ell_{1}$-norm regularization on $\beta$ for all experiments. For the Phi-2 experiment in Section 4.3, we find that $\ell_{2}$-norm works better.
A.2 Forward Pass Only / Semi-structured pruning Experiments
Table 11 shows the Bonsai hyperparameters we used for the experiments in Section 4.1.
@@ A.2 Forward Pass Only / Semi-structured pruning Experiments
A.3 Experiments comparing to Gradient based structured pruning
@@
All Bonsai hyper-parameters are the same as Appendix A.2 except for $\mathrm{ns}_{\mathrm{sub-models}}$ which we set to $1000$.
A.4 Phi-2 pruning experiment details
For the experiment in Section 4.3, All Bonsai hyper-parameters are the same as Appendix A.2 except for the following changes:
@@ A.4 Phi-2 pruning experiment details
Appendix B Impact of regression and perturbation ablation details
For the experiment in Appendix I, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
A simple alternative to Bonsai is to leverage the prior $\rho$, computed from the unperturbed parent model, and make pruning decisions exclusively according to this. This is the No Perturbation + No Regression baseline in Figure 3. This approach has quite poor performance. We can further improve this baseline by adding back perturbative aspect where we prune the parent model according to $\rho^{\prime}$ which is computed by aggregating the $\rho$ metric computed over the perturbed models. Note that we use a Wanda based metric to define $\rho$ for this experiment. module-level analogues of the unstructured pruning metrics we explore are defined in Appendix E.
@@ Appendix B Impact of regression and perturbation ablation details
Linear Regression Relative Speedup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
@@
Appendix C Varying the pruning fraction per-iteration
For the experiment in Appendix I, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $p_{\mathrm{iter}}$.
@@ Appendix C Varying the pruning fraction per-iteration
Prune Frac Relative Speedup w/o Post-Pruning Adaptation w Post-Pruning Adaptation
@@
Appendix D Varying the number of calibration data points for pruning & PPA
For these two categories of experiments, all Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning in the former.
Table 14: How many data points to consider during forward passes. Wikitext-2 Perplexity. Llama-2 7B pruned to 50% sparsity
@@ Appendix D Varying the number of calibration data points for pruning & PPA
Appendix E Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$.
E.1 $\rho$ is Activation Magnitude
@@
Self-Attention Module: For any data-sample sequence $b$, the output of the self-attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho$.
@@
E.2 $\rho$ is Wanda (Sun et al., 2023)
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP. For any data-sample sequence $b$, we flatten model activation before the final output, $\mathbf{a}\in\mathbb{R}^{B\times S\times d}\rightarrow\mathbb{R}^{BS\times d}$ and then compute the following metric which is a module-level analogue of Wanda:
@@
Self-Attention Module: Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the self-attention module. For any data-sample sequence $b$, the output of the self-attention module before the final output projection is $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}$ where $h$ is the number of attention heads and $d_{h}$ is the size of each head’s output. We can flatten $\mathbf{a}\in\mathbb{R}^{B\times S\times d_{h}\times h}\rightarrow\mathbb{R}^{BSd_{h}\times h}$ and then use the same formula as Equation 5 above to calculate $\rho\in\mathbb{R}^{h}$.
E.3 $\rho$ is Fluctuation-based (FLAP) (An et al., 2024)
@@
[…]
where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
@@
[…]
@@
Appendix F How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.